# No time to train! Training-Free Reference-Based Instance Segmentation

### v2

@@ 3.3 Technical implementation details
Our codebase builds upon SAM2-L (Hierarchical ViT) for mask generation and DINOv2-L as the feature encoder. The encoder processes images at 518×518 resolution with a patch size of 14×14, while SAM2 operates at 1024×1024 resolution. During inference, SAM2 first generates candidate masks using a 32×32 grid of query points. For each mask, we compute L2-normalised features by average pooling the encoder features within the masked region. These features are compared against our memory bank, which stores features from $n$ reference images per category. We employ non-maximum suppression with an IoU threshold of 0.5, followed by our semantic-aware soft merging strategy to handle overlapping predictions. The model outputs up to 100 instances per image. The implementation uses PyTorch [1] and PyTorch Lightning [11] for distributed workload across GPUs.
@@ 4.1 Object Detection and Instance Segmentation
COCO-FSOD Benchmark
We evaluate our method in a strict few-shot setting on the COCO-20i dataset [42, 30], using the standard 10-shot [^2] https://paperswithcode.com/sota/few-shot-object-detection-on-ms-coco-10-shot and 30-shot [^3] https://paperswithcode.com/sota/few-shot-object-detection-on-ms-coco-30-shot settings. Results for the COCO-FSOD benchmark are shown in Table 1. All results are reported for COCO-NOVEL classes. Novel classes are the COCO categories that intersect with PASCAL VOC categories [10].
Our method achieves state-of-the-art while being completely training-free, outperforming approaches that fine-tune on novel classes.
Figure 3 presents qualitative results, showing our method’s ability to handle multiple overlapping instances in crowded scenes with fine-grained semantics and precise localisation. With semantic-aware soft merging, we mitigate duplicate detections and false positives. Failure cases are discussed in the Supplementary Material.
Method Ft. on novel 10-shot 30-shot nAP nAP50 nAP75 nAP nAP50 nAP75 TFA [72] 10.0 19.2 9.2 13.5 24.9 13.2 FSCE [66] 11.9 10.5 16.4 16.2 Retentive RCNN [13] 10.5 19.5 9.3 13.8 22.9 13.8 HeteroGraph [18] 11.6 23.9 9.8 16.5 31.9 15.5 Meta F. R-CNN [19] 12.7 25.7 10.8 16.6 31.8 15.8 LVC [32] 19.0 34.1 19.0 26.8 45.8 27.5 C. Transformer [20] 17.1 30.2 17.0 21.4 35.5 22.1 NIFF [16] 18.8 20.9 DiGeo [47] 10.3 18.7 9.9 14.2 26.2 14.8 CD-ViTO (ViT-L) [14] 35.3 54.9 37.2 35.9 54.5 38.0 FSRW [31] 5.6 12.3 4.6 9.1 19.0 7.6 Meta R-CNN [77] 6.1 19.1 6.6 9.9 25.3 10.8 DE-ViT (ViT-L) [82] 34.0 53.0 37.0 34.0 52.9 37.2 Training-free (ours) 36.6 54.1 38.3 36.8 54.5 38.7
Table 1: Comparison of our training-free method against state-of-the-art approaches on the COCO-FSOD benchmark under 10-shot and 30-shot settings. Our approach achieves state-of-the-art performance without finetuning on novel classes (Ft. on novel). Results are reported in terms of nAP, nAP50, and nAP75. nAP refers to mAP for novel classes. Competing methods results are sourced from [14]. Since we are the only method that provides both bounding box and segmentation results, for simplicity we omit segmentation AP on this table.
Method Ft. on novel Novel Split 1 Novel Split 2 Novel Split 3 Avg
@@
FsDetView [75] 25.4 20.4 37.4 36.1 42.3 22.9 21.7 22.6 25.6 29.2 32.4 19.0 29.8 33.2 39.8 29.2
TFA [72] 39.8 36.1 44.7 55.7 56.0 23.5 26.9 34.1 35.1 39.1 30.8 34.8 42.8 49.5 49.8 39.9
Retentive RCNN [13] 42.4 45.8 45.9 53.7 56.1 21.7 27.8 35.2 37.0 40.3 30.2 37.6 43.0 49.7 50.1 41.1
DiGeo [47] 37.9 39.4 48.5 58.6 61.5 26.6 28.9 41.9 42.1 49.1 30.4 40.1 46.9 52.7 54.7 44.0
HeteroGraph [18] 42.4 51.9 55.7 62.6 63.4 25.9 37.8 46.6 48.9 51.1 35.2 42.9 47.8 54.8 53.5 48.0
Meta Faster R-CNN [19] 43.0 54.5 60.6 66.1 65.4 27.7 35.5 46.1 47.8 51.4 40.6 46.4 53.4 59.9 58.6 50.5
CrossTransformer [20] 49.9 57.1 57.9 63.2 67.1 27.6 34.5 43.7 49.2 51.2 39.5 54.7 52.3 57.0 58.7 50.9
LVC [32] 54.5 53.2 58.8 63.2 65.7 32.8 29.2 50.7 49.8 50.6 48.4 52.7 55.0 59.6 59.6 52.3
NIFF [16] (*) 62.8 67.2 68.0 70.3 68.8 38.4 42.9 54.0 56.4 54.0 56.4 62.1 61.2 64.1 63.9 59.4
Multi-Relation Det [12] 37.8 43.6 51.6 56.5 58.6 22.5 30.6 40.7 43.1 47.6 31.0 37.9 43.7 51.3 49.8 43.1
DE-ViT (ViT-S/14) [82] 47.5 64.5 57.0 68.5 67.3 43.1 34.1 49.7 56.7 60.8 52.5 62.1 60.7 61.4 64.5 56.7
DE-ViT (ViT-B/14) [82] 56.9 61.8 68.0 73.9 72.8 45.3 47.3 58.2 59.8 60.6 58.6 62.3 62.7 64.6 67.8 61.4
DE-ViT (ViT-L/14) [82] 55.4 56.1 68.1 70.9 71.9 43.0 39.3 58.1 61.6 63.1 58.2 64.0 61.3 64.2 67.3 60.2
@@
Table 2: AP50 results on the novel classes of the Pascal VOC few-shot benchmark. Competing method results are sourced from [82]. State-of-the-art results are highlighted in bold. (*) indicates that the corresponding implementation is not publicly accessible. Our proposed training-free approach consistently achieves superior performance across all splits, outperforming fine-tuned methods.
@@
[…]
Methods Ft. on novel 1-shot 5-shot $\mathbf{20^{0}}$ $\mathbf{20^{1}}$ $\mathbf{20^{2}}$ $\mathbf{20^{3}}$ mean $\mathbf{20^{0}}$ $\mathbf{20^{1}}$ $\mathbf{20^{2}}$ $\mathbf{20^{3}}$ mean DiffewS [87] 47.7 56.4 51.9 48.7 51.2 52.0 63.0 54.5 54.3 56.0 DiffewS-n [87] 47.1 56.6 53.8 48.3 52.2 57.3 66.5 60.3 58.8 60.7 HSNet [51] 37.2 44.1 42.4 41.3 41.2 45.9 53.0 51.8 47.1 49.5 CyCTR [80] 38.9 43.0 39.6 39.8 40.3 41.1 48.9 45.2 47.0 45.6 VAT [23] 39.0 43.8 42.6 39.7 41.3 44.1 51.1 50.2 46.1 47.9 BAM [35] 43.4 50.6 47.5 43.4 46.2 49.3 54.2 51.6 49.6 51.2 DCAMA [63] 49.5 52.7 52.8 48.7 50.9 55.4 60.3 59.9 57.5 58.3 HDMNet [58] 43.8 55.3 51.6 49.4 50.0 50.6 61.6 55.7 56.0 56.0 Training-free 32.1 46.6 50.9 48.7 44.6 55.2 49.5 60.7 45.5 52.7
Table 4: Performance comparison of strict few-shot semantic segmentation settings (1 and 5-shot) on COCO-$20^{i}$. We aggregate instance-level predictions to allow comparison with semantic segmentation works. Previous methods results are sourced from [87].
PASCAL VOC Few-Shot Benchmark
The PASCAL-VOC dataset [10] consists of 20 classes. For few-shot evaluation, we adopt the standard approach [16], splitting the classes into three groups, each with 15 base and 5 novel classes. As in prior work [82], we report AP50 results on the novel classes.
Table 2 shows that our method outperforms all previous approaches across all splits, achieving state-of-the-art performance across all splits. This holds for both methods that fine-tune on novel classes and those that do not.
4.2 Cross-Domain Few-Shot Object Detection
@@
The CD-FSOD benchmark [14] is designed to evaluate cross-domain few-shot object detection (CD-FSOD) models by addressing challenges in domain shifts and limited data scenarios. It uses COCO [42] as the source training dataset (SD), and six target datasets (TD) — ArTaxOr, Clipart1k, DIOR, DeepFish, NEUDET, and UODD — spanning photorealistic, cartoon, aerial, underwater, and industrial domains with high inter-class variance.
While many approaches fine-tune on a few labeled instances (support set S) from TD before testing on the query set Q, our model is entirely training-free. Thus, we directly evaluate on the six target datasets without any fine-tuning.
Table 3 compares FSOD methods on the CD-FSOD benchmark across 1-shot, 5-shot, and 10-shot settings. Our method sets a new state-of-the-art among training-free approaches and remains competitive with fine-tuned models. These results demonstrate its strong cross-domain generalisation and robustness without requiring retraining.
4.3 Few-shot semantic Segmentation on COCO
Although our method is designed for instance segmentation, we also evaluate it on the COCO-20i Few-Shot Semantic Segmentation benchmark [52]. The 80 COCO classes are divided into four folds [25, 71, 52], each containing 60 base and 20 novel classes. We assess performance on the 20 novel classes under strict 1-shot and 5-shot settings [87]. Results are shown in Table 4.
To adapt our instance segmentation predictions to semantic segmentation, we aggregate all instances of the same class into semantic maps, enabling direct comparison with prior methods. Despite being entirely training-free, our method achieves competitive performance against fine-tuned approaches.
4.4 Variance in reference set
Figure 4: Performance variance across different n-shot settings on the COCO-20i benchmark. We report the mean average precision (mAP) with error bars representing the std over 10 runs with different randomly selected reference images. Results show that variance is higher for lower n-shot settings (e.g., 1–3 shots) due to the greater dependence on reference image selection. As the number of shots increases, the variance decreases, demonstrating the robustness of our training-free method to reference set variations.
Using different reference images leads to variations in results, as performance depends on the quality of the selected reference images for each class.
To quantify this variance, we evaluate our method on the COCO-20i few-shot object detection benchmark using different random seeds to select reference images. Figure 4 shows the standard deviation (std) across 10 runs. We observe that increasing the number of reference images (higher n-shots) reduces result variance, with lower std values. In 1, 2, and 3-shot settings, reference image selection has a more noticeable impact, while for 5 or more shots, the low std demonstrates robustness to reference set variation.
These findings suggest that some reference images are inherently stronger for a given shot setting. Exploring the characteristics of an “optimal” reference set remains an open direction for future work.
5 Conclusions

### v3

@@ 3.3 Technical implementation details
Our code (available in Supp.) builds upon SAM2-L (Hierarchical ViT) for mask generation and DINOv2-L as feature encoder. The encoder processes images at 518×518 resolution with a patch size of 14×14, while SAM2 operates at 1024×1024 resolution. During inference, SAM2 generates candidate masks using a 32×32 grid of query points. For each mask, we compute L2-normalised features by average pooling the encoder features within the masked region. These features are compared against our memory bank, which stores features from $n$ reference images per category. We employ non-maximum suppression with an IoU threshold of 0.5, followed by our semantic-aware soft merging strategy to handle overlapping predictions. The model outputs up to 100 instances per image. The implementation uses PyTorch (Ansel et al., 2024) and PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019) for distributed workload across GPUs.
@@ 4.1 Object Detection and Instance Segmentation
COCO-FSOD Benchmark. We evaluate our method in a strict few-shot setting on the COCO-20i dataset (Lin et al., 2014; Kang et al., 2019a), using the standard 10-shot and 30-shot settings. Results for the COCO-FSOD benchmark are shown in Table 1. All results are reported for COCO-NOVEL classes. Novel classes are the COCO categories that intersect with PASCAL VOC categories (Everingham et al., 2010). Our method achieves state-of-the-art while being completely training-free, outperforming approaches that fine-tune on novel classes. Figure 3 presents qualitative results, showing our method’s ability to handle multiple overlapping instances in crowded scenes with fine-grained semantics and precise localisation. With semantic-aware soft merging, we mitigate duplicate detections and false positives. Failure cases are discussed in the Appendix A.9.
Method FT 10-shot 30-shot nAP nAP50 nAP75 nAP nAP50 nAP75 TFA (Wang et al., 2020) 10.0 19.2 9.2 13.5 24.9 13.2 FSCE (Sun et al., 2021) 11.9 10.5 16.4 16.2 Retentive RCNN (Fan et al., 2021) 10.5 19.5 9.3 13.8 22.9 13.8 HeteroGraph (Han et al., 2021) 11.6 23.9 9.8 16.5 31.9 15.5 Meta F. R-CNN (Han et al., 2022a) 12.7 25.7 10.8 16.6 31.8 15.8 LVC (Kaul et al., 2022) 19.0 34.1 19.0 26.8 45.8 27.5 C. Transformer (Han et al., 2022b) 17.1 30.2 17.0 21.4 35.5 22.1 NIFF (Guirguis et al., 2023) 18.8 20.9 DiGeo (Ma et al., 2023) 10.3 18.7 9.9 14.2 26.2 14.8 CD-ViTO (ViT-L) (Fu et al., 2025) 35.3 54.9 37.2 35.9 54.5 38.0 FSRW (Kang et al., 2019b) 5.6 12.3 4.6 9.1 19.0 7.6 Meta R-CNN (Yan et al., 2019) 6.1 19.1 6.6 9.9 25.3 10.8 DE-ViT (L) (Zhang et al., 2024b) 34.0 53.0 37.0 34.0 52.9 37.2 Training-free (ours) 36.6 54.1 38.3 36.8 54.5 38.7
Table 1: Comparison of our training-free method against state-of-the-art approaches on the COCO-FSOD (10-shot and 30-shot). We achieve SOTA performance without finetuning on novel classes. Results are reported as nAP, nAP50, and nAP75. nAP refers to mAP for novel classes. Competing results are sourced from (Fu et al., 2025). Since we are the only method providing both bounding box and segmentation results, we omit segmentation AP on this table. FT refers to finetuning on novel classes.
PASCAL VOC Few-Shot Benchmark. The PASCAL-VOC dataset (Everingham et al., 2010) consists of 20 classes. For few-shot evaluation, we adopt the standard approach (Guirguis et al., 2023), splitting the classes into three groups, each with 15 base and 5 novel classes. As in prior work (Zhang et al., 2024b), we report AP50 results on the novel classes. Table 2 shows that our method outperforms all previous approaches across all splits, achieving state-of-the-art performance across all splits. This holds for methods that fine-tune on novel classes and those that do not.
Method FT Novel Split 1 Novel Split 2 Novel Split 3 Avg
@@
FsDetView (Xiao and Marlet, 2020) 25.4 20.4 37.4 36.1 42.3 22.9 21.7 22.6 25.6 29.2 32.4 19.0 29.8 33.2 39.8 29.2
TFA (Wang et al., 2020) 39.8 36.1 44.7 55.7 56.0 23.5 26.9 34.1 35.1 39.1 30.8 34.8 42.8 49.5 49.8 39.9
Retentive RCNN (Fan et al., 2021) 42.4 45.8 45.9 53.7 56.1 21.7 27.8 35.2 37.0 40.3 30.2 37.6 43.0 49.7 50.1 41.1
DiGeo (Ma et al., 2023) 37.9 39.4 48.5 58.6 61.5 26.6 28.9 41.9 42.1 49.1 30.4 40.1 46.9 52.7 54.7 44.0
HeteroGraph (Han et al., 2021) 42.4 51.9 55.7 62.6 63.4 25.9 37.8 46.6 48.9 51.1 35.2 42.9 47.8 54.8 53.5 48.0
Meta Faster R-CNN (Han et al., 2022a) 43.0 54.5 60.6 66.1 65.4 27.7 35.5 46.1 47.8 51.4 40.6 46.4 53.4 59.9 58.6 50.5
CrossTransformer (Han et al., 2022b) 49.9 57.1 57.9 63.2 67.1 27.6 34.5 43.7 49.2 51.2 39.5 54.7 52.3 57.0 58.7 50.9
LVC (Kaul et al., 2022) 54.5 53.2 58.8 63.2 65.7 32.8 29.2 50.7 49.8 50.6 48.4 52.7 55.0 59.6 59.6 52.3
NIFF (Guirguis et al., 2023) (*) 62.8 67.2 68.0 70.3 68.8 38.4 42.9 54.0 56.4 54.0 56.4 62.1 61.2 64.1 63.9 59.4
Multi-Relation Det (Fan et al., 2020) 37.8 43.6 51.6 56.5 58.6 22.5 30.6 40.7 43.1 47.6 31.0 37.9 43.7 51.3 49.8 43.1
DE-ViT (ViT-S/14) (Zhang et al., 2024b) 47.5 64.5 57.0 68.5 67.3 43.1 34.1 49.7 56.7 60.8 52.5 62.1 60.7 61.4 64.5 56.7
DE-ViT (ViT-B/14) (Zhang et al., 2024b) 56.9 61.8 68.0 73.9 72.8 45.3 47.3 58.2 59.8 60.6 58.6 62.3 62.7 64.6 67.8 61.4
DE-ViT (ViT-L/14) (Zhang et al., 2024b) 55.4 56.1 68.1 70.9 71.9 43.0 39.3 58.1 61.6 63.1 58.2 64.0 61.3 64.2 67.3 60.2
@@
Table 2: AP50 results on the novel classes of the Pascal VOC few-shot benchmark. Competing method results are sourced from (Zhang et al., 2024b). State-of-the-art results are highlighted in bold. (*) indicates that the corresponding implementation is not publicly accessible. Our proposed training-free approach consistently achieves superior performance across all splits, outperforming fine-tuned methods. FT refers to finetuning on novel classes.
4.2 Cross-Domain Few-Shot Object Detection
@@
Table 3: Per-cell values are 1-shot/5-shot/10-shot mAP on the CD-FSOD benchmark. The $\circ$ symbol indicates results sourced from Distill-cdfsod (Xiong, 2023), while $\dagger$} denotes results reported by CD-ViTO (Fu et al., 2025). ‘Avg.’ represents the average performance across datasets. FT refers to finetuning on novel classes.
@@
The CD-FSOD benchmark (Fu et al., 2025) evaluates cross-domain few-shot object detection models by addressing challenges in domain shifts and limited data scenarios. It uses COCO as the source training dataset (SD), and six target datasets (TD) — ArTaxOr, Clipart1k, DIOR, DeepFish, NEUDET, and UODD — spanning photorealistic, cartoon, aerial, underwater, and industrial domains with high inter-class variance. While many approaches fine-tune on labeled instances (support set S) from TD before testing on the query set Q, our model is training-free. Thus, we directly evaluate on the six target datasets without fine-tuning. Table 3 compares FSOD methods on CD-FSOD across 1, 5, and 10-shot settings. Our method is state-of-the-art among training-free approaches and remains competitive with fine-tuned models. These results demonstrate its strong cross-domain generalisation and robustness without retraining.
Figure 4: Feature comparison across semantic backbones. Left: full image and cropped region of interest. Right: feature PCA maps across backbones. CLIP features are low-resolution and irregular, PE-Spatial is noisy but informative, and DINOv2/DINOv3 are spatially consistent and structured. More visualisations in Figure 19.
Figure 5: Single vs. aggregated feature similarity. We compare cosine similarity maps obtained from (top) a single DINOv2 patch feature selected inside the reference mask (marked with +++++++++++++++++) and (bottom) the aggregated prototype obtained by averaging all features within the mask. For each case, we show intra-class similarity (within the same image) and inter-class similarity (with a target image). Single-feature similarity highlights only local object parts, whereas aggregated features produce more coherent, object-level similarity patterns.
Figure 6: Memory-bank visualisation and model predictions across backbones. (Left) Memory-bank: reference images overlaid with their masked features. (Right) Model predictions: outputs from our pipeline with different backbones (DINOv2, DINOv3, CLIP, and PE-Spatial) along with ground truth masks for comparison. These examples illustrate how differences in feature quality influence downstream detection and segmentation. Additional outputs are provided in Figure 20.
4.3 Ablations
Model Backbone. To evaluate the method transferability across foundation models, we replace the semantic encoder DINOv2 with CLIP (Radford et al., 2021), DINOv3 (Siméoni et al., 2025), and PE-Spatial (Bolya et al., 2025); and the SAM-based segmenter with SAM variants. Results are reported in Table 4. Our pipeline transfers across backbones with moderate performance variation. CLIP models exhibit a performance drop, due to their low-resolution and noisier feature maps. In contrast, PE-Spatial and DINOv3 achieve competitive results with modest degradation. All models are used off-the-shelf without hyperparameter adjustments.
Figure 5 visualises backbones feature maps with PCA. CLIP features show low spatial detail and irregular patterns, while PE-Spatial features are noisy but discriminative. DINOv2 and DINOv3 produce consistent, high-resolution features, aligning with their stronger performance. Figure 6 illustrates the predictions across backbones, alongside memory-bank exemplars with their PCA feature visualisations. More examples are provided in Figure 19 and Figure 20 respectively. These results show that it is not tied to specific backbone and transfers across a range of foundation encoders.
Feature matching. We compare patch-level and prototype-level feature matching using DINOv2 (Fig. 5). While single patch features capture local appearance, they lack full object coverage and yield unstable similarity responses. Averaging features over the reference mask produces object-aligned prototypes, resulting in more spatially coherent and stable similarity maps for memory-bank matching. Additional examples are shown in Figures 21 and 22.
Semantic backbone Segm. backb. 10-shot 30-shot bbox segm bbox segm nAP nAP nAP nAP DINOv2-ViT-L-14 SAM2-L 35.7 33.3 36.8 34.2 DINOv3-ViT-B-16 SAM2-L 32.8 30.9 33.5 31.6 DINOv3-ViT-L-16 SAM2-L 33.8 32.2 34.4 32.8 DINOv3-ViT-H-16+ SAM2-L 25.5 24.4 26.2 25.0 DINOv3-ViT-H-16+ (px) SAM2-L 26.8 25.5 27.6 26.3 CLIP-ViT-B-32 SAM2-L 15.2 13.9 15.8 14.6 CLIP-ViT-B-16 SAM2-L 19.2 18.1 19.5 18.3 CLIP-ViT-L-14 SAM2-L 18.6 17.3 18.7 17.4 CLIP-ViT-L-14-336px SAM2-L 17.8 16.7 17.7 16.7 PE-Spatial-L-14-448 (PE) SAM2-L 25.7 24.1 26.5 24.7 PE-Spatial-L-14-448 (IN) SAM2-L 26.5 24.7 27.3 25.4 PE-Spatial-G-14-448 (PE) SAM2-L 24.7 23.1 24.9 23.3 DINOv2-ViT-L-14 SAM2-T 27.0 26.1 27.9 26.9 DINOv2-ViT-L-14 SAM2-S 29.8 27.7 30.6 28.5 DINOv2-ViT-L-14 SAM2-B+ 29.6 28.4 30.4 29.2
Table 4: Backbone ablation on COCO-FSOD. We compare different semantic encoders (DINOv2, DINOv3, CLIP, PE-Spatial) paired with the same SAM-based (hiera) segmenter. Similarly, we compare different SAM model sizes paired with the same semantic backbone. Results are reported for 10-shot and 30-shot settings using nAP, nAP50, and nAP75. All experiments use a single seed (33) and no hyperparameter tuning (default prompting and aggregation settings). Results show that the pipeline remains functional across encoder choices, with performance largely influenced by feature-map resolution. PE refers to model’s default normalisation, and IN indicates ImageNet normalisation. Model sizes are tiny (T), base (B), large (L), huge (H). Image pixels (px) indicates a larger input image size. 14, 16, 32 refer to the model’s patch size.
Do VLMs provide a competitive alternative? We evaluate a Qwen2.5-VL-7B + SAM2 pipeline on COCO few-shot and observe lower detection and segmentation performance. Details are in Appendix A.3.
Do reference images impact performance? Reference images affect performance in low-shot settings, with variance decreasing as the number of references grows. Analysis of COCO annotations shows that larger, more centered masks yield stronger reference prototypes, motivating simple selection heuristics that improve one-shot performance without adding more references. Our method is robust to substantial reference degradation (e.g., blur), indicating limited sensitivity to reference quality. Full analysis in Appendix A.1.
Pipeline design. We analise pipeline components in Appendix A.4.
Why do similar classes remain challenging across backbones? Semantically similar categories exhibit overlapping feature embeddings, indicating that errors stem from backbone feature geometry rather than prototype construction. DINOv3 shows similar entanglement and produces smoother features that further reduce discriminability during prototype matching (Appendix A.5).
Runtime and memory efficiency. We report inference speed, memory usage, and model size, showing that the training-free pipeline remains efficient with cached references and lightweight matching (Appendix A.6).
5 Conclusion